# A non-abelian version of Deligne’s Fixed Part Theorem

### v4

@@ Appendix B Extensions of representations of profinite groups (after Simpson)
Let $E$ be a finite field extension of ${\mathbb{Q}}_{\ell}$. Let $H\subset G$ be a closed normal subgroup of a profinite group $G$. Set $\Gamma=G/H$. Let $\rho_{H}\colon H\to\mathrm{GL}_{r}(E)$ be a continuous representation and denote by $[\rho_{H}]$ the isomorphism class of $\rho_{H}$. Then $\Gamma$ acts on the set of isomorphism classes of continuous representations $H\to\mathrm{GL}_{r}(E)$ by conjugation.
@@
If $\rho_{H}$ is semi-simple then the stabilizer group $\Gamma_{[\rho_{H}]}$ is closed in $\Gamma$.
@@
If $\rho_{H}$ is absolutely simple, $\det(\rho_{H})$ is finite and $\Gamma$ fixes $[\rho_{H}]$, there essentially exists a unique extension of $\rho_{H}$ to a continuous representation $\rho\colon G\to\mathrm{GL}_{r}(E)$ with finite determinant.
@@
In order to prove essential uniqueness in Proposition B.2 let $\rho$ and $\tilde{\rho}$ be two lifts of $\mathbb{P}\rho$ to two continuous representations of $G$ with finite determinants and which agree on $H$. Then $\tau(g)=\rho(g)\tilde{\rho}(g)^{-1}\in K^{\times}$ defines a continuous homomorphism $\tau\colon\Gamma\to K^{\times}$ with finite image. We deduce $\rho=\tilde{\rho}$ after restricting to the kernel of $\tau$. This proves the unicity part. See Proposition 4.6 for a similar method.
For the existence part in Proposition B.2 we first check that it is sufficient to construct $\rho$ with values in ${\mathrm{GL}}_{r}(\tilde{E})$, where $\tilde{E}$ is a finite Galois extension of $E$. Indeed, then for any $\eta\in\mathrm{Gal}(\tilde{E}/E)$ the extensions $\eta\cdot\rho$ and $\rho$ are essentially equal by the uniqueness part. So, as $\mathrm{Gal}(\tilde{E}/E)$ is finite, we can replace $G$ by an open subgroup containing $H$ so they are equal for all $\eta$, i.e. $\rho$ takes values in ${\mathrm{GL}}_{r}(E)$.
@@
As the restriction of $\mathrm{Ob}(\mathbb{P}\rho)$ to $H$ vanishes, the Hochschild-Serre spectral sequence implies that ${\rm Ob}({\mathbb{P}}\rho)$ lies in a subgroup of $H^{2}_{\rm cont}(G,\mu_{r}(E))$ involving $H^{a}(\Gamma,H^{2-a}(H,\mu_{r}(E)))$ for $a=1,2$ only. Those two groups die on a finite index subgrouop of $\Gamma$. So we can assume that $\mathrm{Ob}(\mathbb{P}\rho)=0$.

### v5

@@ Appendix B Extensions of representations of profinite groups (after Simpson)
Let $E$ be a finite field extension of $\mathbb{Q}$$_{\ell}$. Let $H\subset G$ be a closed normal subgroup of a profinite group $G$. Set $\Gamma=G/H$. Let $\rho_{H}\colon H\to\mathrm{GL}_{r}(E)$ be a continuous representation and denote by $\rho_{H}]$ the isomorphism class of $\rho_{H}$. Then $\Gamma$ acts on the set of isomorphism classes of continuous representations $H\to\mathrm{GL}_{r}(E)$ by conjugation.
@@
If $\rho_{H}$ is semisimple then the stabilizer group $\Gamma_{[\rho_{H}]}$ is closed in $\Gamma$.
@@
If $\rho_{H}$ is absolutely simple, $\det(\rho_{H})$ is finite and $\Gamma$ fixes $\rho_{H}]$, there essentially exists a unique extension of $\rho_{H}$ to a continuous representation $\rho\colon G\to\mathrm{GL}_{r}(E)$ with finite determinant.
@@
In order to prove essential uniqueness in Proposition B.2 let $\rho$ and $\tilde{\rho}$ be two lifts of $\mathbb{P}\rho$ to two continuous representations of $G$ with finite determinants and which agree on $H$. Then $\tau(g)=\rho(g)\tilde{\rho}(g)^{-1}\in K^{\times}$ defines a continuous homomorphism $\tau\colon\Gamma\to K^{\times}$ with finite image. We deduce $\rho=\tilde{\rho}$ after restricting to the kernel of $\tau$. This proves the uniqueness part. See Proposition 4.6 for a similar method.
[…]
@@
As the restriction of $\mathrm{Ob}(\mathbb{P}\rho)$ to $H$ vanishes, the Hochschild-Serre spectral sequence implies that ${\rm Ob}({\mathbb{P}}\rho)$ lies in a subgroup of $H^{2}_{\rm cont}(G,\mu_{r}(E))$ involving $H^{a}(\Gamma,H^{2-a}(H,\mu_{r}(E)))$ for $a=1,2$ only. Those two groups die on a finite index subgroup of $\Gamma$. So we can assume that $\mathrm{Ob}(\mathbb{P}\rho)=0$.